Find the sum: (a) (–7) + (–9) + 4 + 16 (b) (37) + (–2) + (–65) + (–8)
 
rbjm jrmheijt netjmn ij tejmkdfnjgdmn jt

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear student

(a) (-7) + (-9) + 4 + 16
= (-7) + (-9) + 4 + (+7) + (+9)
= (-7) + (+7) + (-9) + (+9) + 4
= 0 + 0 + 4 [(-a) + (+a) = 0]
= 4

(b) (37) + (-2) + (-65) + (-8)
= (+37) + (-75)
= (+37) + (-37) + (-38)
= 0 + (-38) [(-a) + (+a) = 0]
= -38


Thanks